# Optimal Conformal Prediction, E-values, Fuzzy Prediction Sets and Subsequent Decisions

### v1

@@ 3.1 Fuzzy confidence sets
where the event $C_{\alpha}^{z^{n}}(z_{n+1})=1$ indicates that $z_{n+1}$ is fully excluded from the confidence set and $C_{\alpha}^{z^{n}}(z_{n+1})=0$ that it is fuzzy included. The coverage guarantee may then be compactly written as
@@
A fuzzy confidence set $\widetilde{C}_{\alpha}^{z^{n}}:\mathcal{Z}\to[0,1]$ generalizes the binary inclusion-exclusion to a degree of exclusion, where $\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})=1$ means a full exclusion, $\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})=0$ a full inclusion, and $0<\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})<1$ a partial exclusion. We say that a fuzzy confidence set is valid if the average degree of exclusion of the true value $Z_{n+1}$ is bounded by $\alpha$:
@@
In the same way that a binary confidence set is equivalent to a single binary test, as expressed in Theorem 1, a fuzzy confidence set is equivalent to a single fuzzy test $\widetilde{\phi}:\mathcal{Z}^{n+1}\to[0,1]$. Indeed, the fuzzy confidence set $\widetilde{C}_{\alpha}^{z^{n}}$ may be recovered by evaluating the fuzzy test $\widetilde{\phi}$ at different $z\in\mathcal{Z}$ in the sample space
$\widetilde{C}_{\alpha}^{z^{n}}(z)=\widetilde{\phi}(z_{1},\dots,z_{n},z),$
and the fuzzy confidence set is valid if this test $\widetilde{\phi}$ is valid. We delay presenting this result to a slightly more general result in Section 3.2.
@@
If desired, a binary confidence set may be retrieved from a fuzzy confidence set through randomization. In particular, we may independently draw $U\sim\textnormal{Unif}[0,1]$, and construct the (randomized) confidence set
@@
which excludes points for which the degree of exclusion exceeds $U$. This randomized confidence set then still satisfies the classical coverage guarantee, where the coverage probability includes the randomization. We generally do not recommend such a randomized procedure, unless a binary confidence set is practically necessary.
@@ 3.2 Inclusion at different levels $\alpha$
A downside of fuzzy confidence sets is that the ‘degree of exclusion at level $\alpha$’ is difficult to interpret. We overcome this by using recent innovations in the e-value literature developed in Koning (2025b), to merge the degree of exclusion into the significance level. Instead of a ‘fuzzy exclusion’, we view a fuzzy confidence set as excluding values $z\in\mathcal{Z}$ at different significance levels.
Without loss of generality, the first step towards this interpretation is to divide everything by the significance level $\alpha$, moving from the $[0,1]$-scale to the $0,1/\alpha]-$scale. We refer to this scale as the evidence scale. To avoid ambiguity, we denote the fuzzy confidence set on the evidence scale by $\mathcal{E}^{z^{n}}:\mathcal{Z}\to[0,1/\alpha]$, defined as
@@
Due to the division by $\alpha$, it is said to be valid if
@@
A consequence of the rescaling is that a full-exclusion of $z$ at significance level $\alpha$ now corresponds to $\mathcal{E}^{z^{n}}(z)=1/\alpha$ for a $0,1/\alpha]-$valued fuzzy confidence set. For example, suppose that we pre-specify $\alpha=0.05$. Then, $\mathcal{E}^{z^{n}}(z)=20$ corresponds to an exclusion of $z$ at level $1/20=0.05$. A natural question is then whether $\mathcal{E}^{z^{n}}(z)=10$, say, may be interpreted as an exclusion of level $\alpha^{\prime}=1/10=0.1$? Remarkably, it turns out that the answer is yes, if we appropriately handle the fact that this makes the significance level data-dependent (Grünwald, 2023a; Koning, 2024). More generally, a fuzzy confidence set $\mathcal{E}^{z^{n}}:\mathcal{Z}\to[0,\infty]$ may be interpreted as reporting for each value $z\in\mathcal{Z}$ the smallest significance level at which it is excluded from the confidence set.
We now formalize precisely what we mean with this statement. To start, we move to a different representation of our fuzzy confidence set $\mathcal{E}^{Z^{n}}$ as a collection of its $\alpha$-labeled sublevel sets $(C_{\alpha}^{Z^{n}})_{\alpha>0}$, by collecting all $z\in\mathcal{Z}$ for which the evidence is less than $1/\alpha$:
$C_{\alpha}^{Z^{n}}=\{z\in\mathcal{Z}:\mathcal{E}^{Z^{n}}(z)<1/\alpha\}.$ (8)
An intermediate observation is that these sets $C_{\alpha}^{Z^{n}}$ are each, individually, valid confidence sets:
$\frac{P(Z_{n+1}\not\in C_{\alpha}^{Z^{n}})}{\alpha}$ $=\frac{P(\mathcal{E}^{Z^{n}}(Z_{n+1})\geq 1/\alpha)}{\alpha}$
@@
where the first inequality follows from Markov’s inequality, and the second inequality follows from the validity of the fuzzy confidence set (7).
Markov’s inequality is generally extremely wasteful. We may close the slack introduced in (9) by Markov’s inequality by instead showing that the family of confidence sets $(C_{\alpha}^{Z^{n}})_{\alpha>0}$ together satisfies a much stronger joint coverage guarantee for data-dependent levels. This may be viewed as a generalization of (9) to data-dependent significance levels, stating that (9) must hold ‘on average’ over the data-dependent level. The proof mimics that of Theorem 2 in Koning (2024), applied to the exclusion event $\mathbb{I}\{Z_{n+1}\not\in C_{\alpha}\}$. For completeness, we include it in Appendix A.1. The subsequent Lemma 1 follows immediately from (8).
@@
The collection of sublevel sets $(C_{\alpha}^{Z^{n}})_{\alpha>0}$ of a valid fuzzy confidence set $\mathcal{E}^{Z^{n}}$ satisfies
$\mathrm{E}_{\widetilde{\alpha}}^{P}\left[\frac{P(Z_{n+1}\not\in C_{\widetilde{\alpha}}^{Z^{n}}\mid\widetilde{\alpha})}{\widetilde{\alpha}}\right]\leq 1,\textnormal{ for every }P\in\mathcal{P},$ (10)
@@
The smallest data-dependent level $\widetilde{\alpha}$ for which $z$ is excluded from $C_{\widetilde{\alpha}}$ equals $\widetilde{\alpha}=1/\mathcal{E}^{Z^{n}}(z)$.
@@
We end with Theorem 4, which generalizes the equivalence between prediction confidence sets and tests from Theorem 1 to fuzzy prediction confidence sets and e-values. Its proof is omitted, as it mimics that of Theorem 1, plugging in e-values for tests and fuzzy confidence sets for confidence sets.
@@
A fuzzy confidence set $\mathcal{E}^{Z^{n}}:\mathcal{Z}\to[0,\infty]$ is equivalent to an e-value $\varepsilon:\mathcal{Z}^{n+1}\to[0,\infty]$, which is valid for $\mathcal{P}$ if and only if the e-value is valid for $\mathcal{P}$.
@@
Predictive distributions (Vovk et al., 2017) may be interpreted as defective fuzzy confidence sets. Indeed, a standard predictive distribution is equivalent to taking a collection of confidence sets $(C_{\alpha}^{Z^{n}})_{\alpha>0}$, which are each merely individually valid at data-independent levels as in (9), and reporting for each $z$ the smallest data-dependent level $\alpha$ at which it is excluded. This means predictive distributions contain a mismatch between their guarantee (for data-independent levels) and what they report (an exclusion at a data-dependent level). A consequence is that if we wish to retain the guarantee in subsequent decisions, then we must pre-specify $\alpha$. In turn, we may only use the corresponding confidence set $C_{\alpha}^{Z^{n}}$ from the collection, which renders the remainder of the confidence distribution redundant.
@@
Fuzzy confidence sets resolve this mismatch, as their validity guarantee (10) is over data-dependent levels. As a consequence, our fuzzy confidence sets may also be interpreted as precisely those kind of predictive distributions that do permit us to look at the entire confidence distribution and still retain a guarantee.
3.3 Optimal fuzzy confidence sets
@@
In this section, we study optimal fuzzy confidence sets, which generalizes the optimality theory for the classical binary confidence sets developed in Section 2.3. The classical binary setting from Section 2.3 is equivalent to choosing the binary e-value $\varepsilon:\mathcal{Z}^{n+1}\to\{0,1/\alpha\}$ that maximizes the objective
$\mathrm{E}^{Q}\left[\varepsilon\right].$ (11)
@@
The naive natural fuzzy generalization would be to choose the $0,1/\alpha]-$valued e-value that maximizes this same objective. This indeed minimizes the expected size of the confidence set, if we were to permit randomization as in Remark 1. Moreover, this corresponds to the classical Neyman–Pearson most powerful randomized test for $\mathcal{P}$ against $Q$, and is the kind of objective considered by Geyer and Meeden (2005) in the context of a classical fuzzy confidence set on a parameter space.
@@
However, to truly unlock the potential of fuzzy confidence sets we consider more general expected-utility objectives (Koning, 2025b). To motivate this, note that the objective (11) over $0,1/\alpha$]-valued e-values is equivalent to maximizing the objective
$\mathrm{E}^{Q}[\varepsilon\wedge 1/\alpha],$
@@
over unbounded e-values $\varepsilon:\mathcal{Z}^{n+1}\to[0,\infty]$, because this objective places no value on evidence beyond $1/\alpha$. This reveals that the objective (11) may be viewed as maximizing the expected utility
$\mathrm{E}^{Q}[U(\varepsilon)],$
under the utility function $U:x\mapsto x\wedge 1/\alpha$. This utility function may be interpreted as expressing the quite rigid preference that we prefer more evidence linearly, up to some threshold $1/\alpha$. That is, we have no interest in seeing exclusions at levels smaller than $\alpha$.
@@
As we do not believe this is the true preference of analysts using prediction confidence sets, we extend to more general utility functions $U:[0,\infty]\to[0,\infty]$. Here, we assume the utility function is non-decreasing and concave, which expresses that we care decreasingly much about obtaining more evidence. In the e-value literature, the most popular utility is log-utility $U:x\mapsto\log(x)$ (Grünwald et al., 2023; Larsson et al., 2024). However, more general utilities have also been studied (Koning, 2025b). One example are power-utilities $U:x\mapsto((x^{h})-1)/h$, $h\neq 0$, where the log-utility appears as the limit as $h\to 1$, or bounded versions of such utilities.
@@
In fact, to obtain properly fuzzy confidence sets, we must move away from Neyman–Pearson-style utility functions $U:x\mapsto x\wedge 1/\alpha$. This is because Neyman–Pearson-style utility function generally produce barely-fuzzy confidence sets, with the fuzziness only appearing in discrete settings. We suspect that Geyer and Meeden (2005) did not realize the potential of fuzzy confidence sets outside of discrete data, as they implicitly stuck to such a Neyman–Pearson-style utility. By extending to more general utilities we extend the scope of fuzzy confidence sets to non-discrete settings.
@@
Let $\mu^{|Z^{n}}$ be a $Z^{n}$-dependent probability kernel, and $U:[0,\infty]\to[0,\infty]$. Then $\mathcal{E}$ maximizes
@@
if and only if $\varepsilon$ is the $U$-expected-utility optimal e-value against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}}$.

### v2

@@ 4.1 From classical to fuzzy prediction sets
where the event $C_{\alpha}^{z^{n}}(z_{n+1})=1$ indicates that $z_{n+1}$ is excluded from the prediction set and $C_{\alpha}^{z^{n}}(z_{n+1})=0$ that it is included. The coverage guarantee may then be compactly written as
@@
A fuzzy prediction set $\widetilde{C}_{\alpha}^{z^{n}}:\mathcal{Z}\to[0,1]$ generalizes the binary inclusion-exclusion to a degree of exclusion, where $\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})=1$ means a full exclusion, $\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})=0$ a full inclusion, and $0<\widetilde{C}_{\alpha}^{z^{n}}(z_{n+1})<1$ a partial exclusion. We say that a fuzzy prediction set is valid if the average degree of exclusion of the true value $Z_{n+1}$ is bounded by $\alpha$:
@@
As captured in Theorem 1 for binary prediction sets, a fuzzy prediction set $\widetilde{C}_{\alpha}^{Z^{n}}$ for $Z_{n+1}$ is a slice of a fuzzy prediction set $\widetilde{C}_{\alpha}:\mathcal{Z}^{n+1}\to[0,1]$ for $Z^{n+1}$:
$\widetilde{C}_{\alpha}^{z^{n}}(z)=\widetilde{C}_{\alpha}(z_{1},\dots,z_{n},z).$
Moreover, the fuzzy prediction set $\widetilde{C}_{\alpha}^{Z^{n}}$ for $Z_{n+1}$ is valid if $\widetilde{C}_{\alpha}$ is valid. We delay presenting this result to Section 4.2, where we present it in a slightly more general form in Theorem 4.
@@
If desired, a binary prediction set may be retrieved from a fuzzy prediction set through randomization. In particular, we may independently draw $U\sim\textnormal{Unif}[0,1]$, and construct the (randomized) prediction set
@@
which excludes points for which the degree of exclusion exceeds $U$. This randomized prediction set then still satisfies the classical coverage guarantee, where the coverage probability includes the randomization. We generally do not recommend such a randomized procedure unless a binary prediction set is practically necessary.
@@ 4.2 Inclusion at different levels $\alpha$
A downside of fuzzy prediction sets is that the ‘degree of exclusion at level $\alpha$’ is difficult to interpret. We overcome this by using recent innovations in the e-value literature developed in koning2025continuoustestingunifyingtests, to merge the degree of exclusion into the significance level. Instead of a ‘fuzzy exclusion’, we view a fuzzy prediction set as excluding values $z\in\mathcal{Z}$ at different significance levels.
Without loss of generality, the first step towards this interpretation is to divide everything by the significance level $\alpha$, moving from the $[0,1]$-scale to the $0,1/\alpha]-$scale. We refer to this scale as the evidence scale. To avoid ambiguity, we denote the fuzzy prediction set on the evidence scale by $\mathcal{E}^{z^{n}}:\mathcal{Z}\to[0,1/\alpha]$, defined as
@@
Due to the division by $\alpha$, it is now said to be valid if
@@
A consequence of this rescaling is that a full-exclusion of $z$ at significance level $\alpha$ now corresponds to $\mathcal{E}^{z^{n}}(z)=1/\alpha$ for a $0,1/\alpha]-$valued fuzzy prediction set. For example, suppose that we pre-specify $\alpha=0.05$. Then, $\mathcal{E}^{z^{n}}(z)=20$ corresponds to an exclusion of $z$ at level $1/20=0.05$. A natural question to ask is whether $\mathcal{E}^{z^{n}}(z)=10$, say, may be interpreted as an exclusion of level $\alpha^{\prime}=1/10=0.1$? Remarkably, it turns out that the answer is yes, if we appropriately handle the fact that this makes the significance level $\alpha^{\prime}$ data-dependent (grunwald2022beyond; koning2024post). More generally, a fuzzy prediction set $\mathcal{E}^{z^{n}}:\mathcal{Z}\to[0,\infty]$ may be interpreted as reporting for each value $z\in\mathcal{Z}$ the smallest significance level at which it is excluded from the prediction set.
We now formalize precisely what we mean with this statement. To start, we move to a different representation of our fuzzy prediction set $\mathcal{E}^{Z^{n}}$ as a collection of its $\alpha$-labeled sublevel sets $(\overline{C}_{\alpha}^{Z^{n}})_{\alpha>0}$, by collecting all $z\in\mathcal{Z}$ for which the evidence is less than $1/\alpha$:
$\overline{C}^{Z^{n}}=\{z\in\mathcal{Z}:\mathcal{E}^{Z^{n}}(z)<1/\alpha\}.$ (8)
An intermediate observation is that these sets $\overline{C}_{\alpha}^{Z^{n}}$ are each, individually, valid prediction sets:
$\frac{P(Z_{n+1}\not\in\overline{C}_{\alpha}^{Z^{n}})}{\alpha}$ $=\frac{P(\mathcal{E}^{Z^{n}}(Z_{n+1})\geq 1/\alpha)}{\alpha}$
@@
where the first inequality follows from Markov’s inequality, and the second inequality follows from the validity of the fuzzy prediction set (7).
Markov’s inequality generally has a lot of slack. In Theorem 3, we close this slack by showing that the family of prediction sets $(\overline{C}_{\alpha}^{Z^{n}})_{\alpha>0}$ together satisfies a much stronger joint coverage guarantee for data-dependent levels. This may be viewed as a generalization of (9) to data-dependent significance levels, stating that (9) must hold ‘on average’ over the data-dependent level.
The proof immediately follows from Theorem 2 in koning2024post, applied to the test $\overline{C}_{\alpha}=\mathbb{I}\{Z_{n+1}\not\in\overline{C}_{\alpha}^{Z^{n}}\}$. For completeness, we include it in Appendix A.1. The subsequent Lemma 1 follows immediately from (8).
@@
The collection of sublevel sets $(\overline{C}_{\alpha}^{Z^{n}})_{\alpha>0}$ of a valid fuzzy prediction set $\mathcal{E}^{Z^{n}}$ satisfies
$\mathrm{E}_{\widetilde{\alpha}}^{P}\left[\frac{P(Z_{n+1}\not\in\overline{C}_{\widetilde{\alpha}}^{Z^{n}}\mid\widetilde{\alpha})}{\widetilde{\alpha}}\right]\leq 1,\textnormal{ for every }P\in\mathcal{P},$ (10)
@@
The smallest data-dependent level $\widetilde{\alpha}$ for which $z$ is excluded from $\overline{C}_{\widetilde{\alpha}}$ equals $\widetilde{\alpha}=1/\mathcal{E}^{Z^{n}}(z)$.
@@
We end with Theorem 4, which generalizes Theorem 1 to fuzzy prediction sets on the evidence $0,\infty]-$scale. This (slightly) generalizes the discussion in Section 4.1, as the discussion there technically does not cover fuzzy prediction sets on an unbounded evidence scale $0,\infty]$.
A fuzzy test on the evidence $0,\infty]-$scale is also known as an e-value (koning2025continuoustestingunifyingtests). Hence, in order to construct a fuzzy prediction set $\mathcal{E}^{Z^{n}}$, it suffices to construct an e-value and then slice it at $Z^{n}$.
@@
A fuzzy prediction set $\mathcal{E}^{Z^{n}}:\mathcal{Z}\to[0,\infty]$ for $Z_{n+1}$ is a slice
$\mathcal{E}^{z^{n}}(z)=\mathcal{E}(z^{n},z)$
of a fuzzy prediction set $\mathcal{E}$ for $Z^{n+1}$. Such a fuzzy prediction set $\mathcal{E}$ is an e-value, and it is valid for $\mathcal{P}$ if and only if it is a valid e-value for $\mathcal{P}$.
@@
Predictive distributions (vovk2017nonparametric) may be interpreted as fuzzy prediction sets with weaker error guarantees. Indeed, a standard predictive distribution is equivalent to taking a collection of prediction sets $(C_{\alpha}^{Z^{n}})_{\alpha>0}$, which are each merely individually valid at data-independent levels as in (9), and reporting for each $z$ the smallest data-dependent level $\alpha$ at which it is excluded. This means predictive distributions contain a mismatch between their guarantee (for data-independent levels) and what they report (an exclusion at a data-dependent level). A consequence is that if we wish to retain the guarantee then we must pre-specify $\alpha$. In turn, we may only use the corresponding prediction set $C_{\alpha}^{Z^{n}}$ from the collection, which renders the remainder of the predictive distribution redundant.
@@
Fuzzy prediction sets resolve this mismatch, as their validity guarantee (10) is over data-dependent levels. As a consequence, our fuzzy prediction sets may also be interpreted as precisely those kind of predictive distributions that do permit us to look at the entire predictive distribution and still retain a guarantee.
4.3 Optimal fuzzy prediction sets
@@
In this section, we study optimal fuzzy prediction sets. This generalizes the optimality theory for the classical binary prediction sets developed in Section 3.3. Instead of viewing this problem as finding an optimal test, we now consider the more general problem of finding an optimal e-value. The classical binary setting from Section 3.3 is equivalent to choosing the binary valid e-value $\mathcal{E}:\mathcal{Z}^{n+1}\to\{0,1/\alpha\}$ that maximizes the objective
$\mathrm{E}^{Q}\left[\mathcal{E}\right].$ (11)
@@
The naive natural fuzzy generalization would be to choose the $0,1/\alpha]-$valued e-value that maximizes this same objective. This indeed minimizes the expected size of the prediction set, if we were to permit randomization as in Remark 3. Moreover, this corresponds to the classical Neyman–Pearson most powerful randomized test for $\mathcal{P}$ against $Q$, and is the kind of objective considered by geyer2005fuzzy in the context of a classical fuzzy prediction set on a parameter space.
@@
However, to truly unlock the potential of fuzzy prediction sets we consider more general expected-utility objectives (koning2025continuoustestingunifyingtests). To motivate this, note that an optimizer of (11) among $0,1/\alpha$]-valued e-values is also an optimizer of the objective
$\mathrm{E}^{Q}[\mathcal{E}\wedge 1/\alpha],$
@@
over unbounded e-values $\mathcal{E}:\mathcal{Z}^{n+1}\to[0,\infty]$, because this objective places no value on evidence beyond $1/\alpha$. This reveals that the objective (11) may be viewed as maximizing the expected utility
$\mathrm{E}^{Q}[U^{\textnormal{NP}}(\mathcal{E})],$
under the utility function $U^{\textnormal{NP}}:x\mapsto x\wedge 1/\alpha$. We name this the “Neyman–Pearson utility function”, as classical Neyman–Pearson hypothesis testing can be viewed as collecting evidence under this utility function, which expresses linear valuation of evidence up to some threshold $1/\alpha$. This threshold expresses that we put no additional value on seeing exclusions at levels smaller than $\alpha$.
@@
As we do not believe this Neyman–Pearson utility function expresses the true preference of analysts using prediction sets, we extend to more general utility functions $U:[0,\infty]\to[-\infty,\infty]$. Here, we assume the utility function is non-decreasing and concave, which expresses that we care decreasingly much about obtaining more evidence. See Remark 5 for a discussion on the choice of utility.
@@
To obtain p8roperly fuzzy prediction sets, we must in fact move away from Neyman–Pearson-style utility functions $U:x\mapsto x\wedge 1/\alpha$. This is because a Neyman–Pearson-style utility function generally produces barely-fuzzy prediction sets, with the fuzziness only appearing in discrete settings. We suspect that geyer2005fuzzy did not appreciate the potential of fuzzy prediction sets outside of discrete data, as they implicitly stuck to such a Neyman–Pearson-style utility. By extending to more general utilities we extend the scope of fuzzy prediction sets to non-discrete settings.
@@
Let $\mu^{|Z^{n}}$ be a $Z^{n}$-dependent probability kernel, and $U:[0,\infty]\to[-\infty,\infty]$. Then $\mathcal{E}^{*}$ maximizes
@@
if and only if $\mathcal{E}$ is the $U$-expected-utility optimal e-value against the alternative $Q=P_{*}^{Z^{n}}\otimes\mu^{|Z^{n}}$.
Remark 5 (Choice of utility function).
In general, the utility $U(e)$ of a realization $e$ of our e-value expresses how much we value a rejection at level $e=1/\alpha$. In the e-value literature, the most popular utility is log-utility $U:x\mapsto\log(x)$ (grunwald2023safe; larsson2024numeraire). The log-utility function may be interpreted as valuing evidence linearly in the order of magnitude: $\log(10)$ is half as valuable as $\log(100)=2\log(10)$. The main motivation for log-utility usually comes from i.i.d. sequential settings, where log-utility optimality can be coupled to minimizing expected stopping times through Wald’s Identity.
In the fuzzy prediction context, such sequential arguments seem less important, so that we see no direct reason to prefer log-utility over other utilities, beyond it serving as a well-studied ‘default’ option. Luckily, more general utilities have also been studied (koning2025continuoustestingunifyingtests). One example are power-utilities $U:x\mapsto(x^{h}-1)/h$, $h\neq 0$, where the log-utility appears as the limit as $h\to 0$, or bounded versions of such utilities.
4.4 Merging fuzzy prediction sets
Fuzzy prediction sets inherit the desirable merging properties of their underlying e-values.
If we have two arbitrary (possibly highly dependent) fuzzy prediction sets $\mathcal{E}_{1}^{Z^{n}}$ and $\mathcal{E}_{2}^{Z^{n}}$ for $Z_{n+1}$, then their average is still a fuzzy prediction set. Indeed, their underlying e-values, say $\mathcal{E}_{1}$ and $\mathcal{E}_{2}$, still average to a valid e-value:
$\overline{\mathcal{E}}=(\mathcal{E}_{1}+\mathcal{E}_{2})/2.$
The resulting prediction set $\overline{\mathcal{E}}^{Z^{n}}$ is then a fuzzy prediction set for $Z_{n+1}$. In fact, this is equivalent to simply averaging the prediction sets $\mathcal{E}_{1}^{Z^{n}}$ and $\mathcal{E}_{2}^{Z^{n}}$ for $Z_{n+1}$. This extends from averaging a finite number of prediction sets to mixtures over collections of prediction sets. wang2025onlyadmissible recently proved that (weighted) averaging (possibly with the constant e-value $\mathcal{E}\equiv 1$) is the only admissible way to merge arbitrarily dependent e-values.
Such arbitrarily dependent fuzzy prediction sets may be different fuzzy prediction sets produced on the same data. For example, this enables us to construct an optimal fuzzy prediction set for many choices of $Q$ and produce a ‘mixture’ fuzzy prediction set that weights our preference over options of $Q$. Or it allows us to average over split-conformal fuzzy prediction sets over different choices of the split (vovk2024conformal).
As non-fuzzy prediction sets (possibly of different levels) can be viewed as special fuzzy prediction sets, they can also be merged in the same manner, but the outcome will generally be fuzzy merged prediction sets. Rounding down such a merged fuzzy prediction set to a non-fuzzy prediction set usually discards a lot of evidence.
Another popular method to merge e-values is through multiplication, which is valid under independence. While averaging fuzzy prediction sets cannot increase evidence beyond the maximum evidence for any of the individual prediction sets, multiplying them does have this potential. At the same time, multiplication may also kill-off evidence if one of the fuzzy prediction sets equals zero at some points. Non-fuzzy prediction sets suffer from this issue, as they have zero-values by construction. The assumption of independence may be dramatically weakened to a kind of conditional mean-independence property of one of the fuzzy prediction sets, given the other: $\mathrm{E}[\mathcal{E}_{2}\mid\mathcal{E}_{1}]\leq 1$.